OpenStudy (anonymous):

Jim pushes on a large bookcase that has a mass of 35kg with a horizontal force of 150N. The coefficient of friction between the wooden floor and bookcase is 0.6. If the bookcase does not move what is the magnitude of the static frictional force (from the floor) acting on the bookcase? Answer in Newtons.

OpenStudy (aaronq):

\(\sum F_x=F_{push}- F_{friction}=0\) \(\rightarrow F_{friction}=\mu F_N\) \(\sum F_y=F_N-F_g=0\rightarrow F_N=F_g=mg \) \(\rightarrow F_{friction}=\mu F_N=\mu mg\)